An empty swimming pool with a square floor dimension has a volume of 3y 2 m3 . Given the weight of air in the swimming pool is about 500 N and the air density is 1.29 kg/m3 . (a) Find the length of the swimming pool’s floor, y. (b) Determine the force exerted by air on the floor of the swimming pool


1
Expert's answer
2021-12-29T13:01:54-0500

Answer

Given data in question

Floor volume V="3y^2 m^3"

Weight of air W=500N

Density "\rho=1.29\frac{kg}{m^3}"

Now

a) length of the swimming pool’s floor, y

The weight can written as

"W=mg"

"=\rho Vg"

"=\rho \times3y^2\times g"

So length of swimming pools floor

"y=\sqrt{\frac{W}{3\rho g}}"

Putting all values

"y=\sqrt{\frac{500}{3\times1.29\times 9.81}}\\y=3.63m"


b) now force exerted by air on the floor of the swimming pool is given by

The equivalent of weight of air

F=W=500N.